### v2

# DFT modelling of stacking faults in hexagonal and cubic GaN

Abstract
We have performed density functional theory (DFT) calculations to characterize the energetics, and the atomic and electronic structure, of stacking faults in GaN, both in the stable hexagonal wurtzite (wz) phase and in the metastable cubic zincblende (zb) phase. In wz GaN, SFs on the (0001) planes can be divided into three different intrinsic stacking faults (I1, I2, and I3) and one extrinsic stacking fault (E). Based on the calculated formation energy, I1 is the most stable SF of wz GaN in agreement with experiment. In zb GaN, SFs form along (111) directions, giving one type each of intrinsic, extrinsic and twin SFs. In our calculations, the three types of SFs have similar formation energy. To characterize the effect of the stacking faults on the electronic structure of the material, we examined the band density. We found that the bands near the valence band maximum in wz GaN are localised on the Ga-polar side of the stacking fault (i.e. on the Ga side of the Ga-N bonds perpendicular to the SF), with the bands near the conduction band minimum more on the N-polar side, though somewhat delocalised. We found the opposite trend in zb GaN; this behaviour is caused by a redistribution of charge near the interface. We also show the band offsets for the stacking faults, finding that they are very sensitive to local conditions, but can all be described as type II interfaces, with the presence of a stacking fault reducing the gap locally in both wz and zb.
I Introduction
III-V nitrides are increasingly important in electronic and optoelectronic devices [13, 19]. These applications are affected by a number of material problems, mostly due to the lack of cost-effective suitable substrates on which the materials are grown. Growth on mismatched substrates causes the epilayers to contain a very high concentration of extended defects, principally dislocations in c-type wz GaN [21, 24] and stacking faults in non-polar oriented wz [17, 42]. Most commercial GaN devices are based on the stable $c$-oriented wurtzite (wz) GaN, which shows a spontaneous polarisation; moreover, in the strained heterostructures required for devices, piezoelectric polarisation is induced [35]. As a result of this phenomenon, band bending occurs through the quantum confined Stark effect (QCSE) [5, 27].
Through epitaxial growth of GaN films on cubic substrates, a metastable cubic zincblende (zb) structure can be stabilized [36, 19], which is not polar like the wz phase, providing a potential improvement in issues associated with the QCSE that occur in wz optoelectronic devices. Recent studies show that, due to the absence of polarization fields and the lower band gap compared to wz GaN, zb GaN may show higher efficiency in longer wavelength emission (green, amber, red) [10]. The stacking faults in zb GaN are often considered as wz insertions in a zb matrix [6] which modify the optical properties of zb GaN films and propagate into active layers (though we show here that this model is not correct). In quantum wells and electron blocking layers, the segregation of alloying elements has been observed at stacking faults, leading to the formation of quantum wires and polarized emission [12].
Many questions have yet to be answered concerning the high resolution characterization of SFs’ structures, alloy segregation to SFs and their influence on the optoelectronic properties of GaN-based devices. Recently, a combination of high-resolution scanning transmission electron microscopy (STEM) and energy dispersive x-ray spectrometry was used to investigate the effects of alloy segregation around stacking faults in a zb GaN light-emitting structure [12]. A key question remaining is to identify the change in atomic spacing at the SFs, and whether this plays a significant role in segregation phenomena, which has been suggested as a possible mechanism for the different behaviour of Al and In [12]: here, atomistic modelling can complement experiments, and provide insights beyond experimental resolution. DFT modelling has previously focused on the energetics and electronic structure of basal stacking faults of wz GaN [32, 3, 4] with forcefield modelling considering energetics in zb GaN [1]. However, there have been no DFT studies on SFs in zb GaN, and no examination of the effects of stacking faults on the local electronic structure and band edges in either wz or zb GaN.
In this paper, our aim is to study the atomic structure and energetics of SFs, alongside their electronic structure and differences to bulk material in both wz and zb phases of GaN, and to compare to recent experimental findings on those SFs (e.g. wz GaN [2, 42, 25] and zb GaN [41, 40]). The rest of the paper is laid out as follows: in Sec. II, the computational approach is presented; the results for the structural and energetic properties are compared in Sec. III; the electronic structure of the SFs is discussed in Sec. IV; Sec. V gives a brief summary and concludes the paper.
II Computational Methods and Models
[…]
The structural relaxation was performed in two steps. First, the simulation cell was optimized to a tolerance of 0.1 GPa using the conjugate gradient algorithm, with fixed fractional ionic coordinates; to mimic the effects of a fixed substrate, only the stress along $c$ was minimised, with values of $a$ and $b$ fixed to bulk parameters. The stabilized quasi-Newton method (SQNM)[30] was then used to optimize the ionic positions with a threshold of 0.02 eVÅ-1.
[…]
The formation energy per unit area of the simulation cell for the optimized SFs is calculated as:
$\Delta E=E_{SF}-nE_{Ref}$ (1)
where $E_{SF}$ is the total energy of the specific SF in a simulation cell containing $n$ layers, and $E_{Ref}$ is the energy per layer of the host material.
When considering band offsets between perfect bulk cells and cells with stacking faults, we need a common zero. The semi-core $d$ electrons in Ga form a natural low-lying band which allows comparison; however there are two groups of low-lying $d$ bands separated by around 1.9 eV, with the lowest energy band showing dispersion of around 0.7 eV, while the upper group of $d$ bands show dispersion of less than 0.1 eV. This upper group forms a suitable reference point to compare the energies of different sets of bands, and we choose the value at the gamma point to set the common zero.
III Physical structure and energetics
Taking the optimized lattice parameters and atomic coordinates of bulk structures, a series of simulations cells was constructed containing stacking faults, with progressively increasing separation between the stacking faults to test the convergence. For wz GaN, the three intrinsic stacking faults I1, I2, I3 in isolation would be formed as …AB/CB…, …AB/CA…and …AB/C/BA…respectively. When modelled with periodic boundary conditions some adjustment is required, and we need two stacking faults in each cell. For n layers of each stacking sequence, we then have /nAB/nCB/, /nAB/nCA/C/ and /nAB/C/nBA/C/ where the stacking faults are in the middle and at the edges of the cell, and an extra C layer has been added to both I2 and I3 to restore periodicity. For the extrinsic stacking fault, we have …AB/C/AB…in isolation, and /nAB/C/ in a periodic cell (in this case, we need only one stacking fault per simulation cell).
For zb GaN, in isolation, the intrinsic, extrinsic and twin stacking faults are formed as …ABC/BC/ABC…, …ABC/B/ABC… and …ABC/BAC/ABC… respectively. We can accommodate a single stacking fault in a periodic cell in all three cases, which for n layers we write as /nABC/BC/, /nABC/B/ and /nABC/BAC/ in each case. These models are illustrated in Figs. 1 and 2.
Figure 1: Stacking faults in wurtzite GaN. (a) Intrinsic-1 (I1) /AB/CBCB/AB/, (b) Intrinsic-2 (I2) /AB/CACA/C/AB/, (c) Intrinsic-3 (I3) /AB/C/BABA/C/AB/ and (d) Extrinsic /ABAB/C/ABAB/. Note that these are highly shortened cells for illustration only.
Figure 2: Stacking faults in Zincblende GaN. (a) Extrinsic /ABC/B/ABC/, (b) Intrinsic /ABC/BC/ABC/ and (c) Twin /ABC/BAC/ABC/. Note that these are highly shortened cells for illustration only.
We tested the effect of increasing the number of layers of bulk material between stacking faults in both wz and zb, to find the converged value of the formation energy, which we give in mJ/m2. The converged values are given in Table 1 along with number of layers. Full details of the convergence of the energy with number of layers are given in the supplementary information. Based on these calculated formation energies, SFs in wz GaN all show a positive energy, with I1 being the most stable SF. This agrees well with the experimental observation that 90% of observed basal plane stacking faults in wz GaN films are of I1-type [42, 26]. The second most common SF is I2 followed by I3 and E; experimentally, the extrinsic type of SF is rarely observed [25]. The observed trend is in good accord with energies calculated in other DFT investigations [32, 3, 4] though none of these studies report details of convergence tests of spacing between SFs.
In zb GaN, however, all stacking faults have a negative formation energy; this is not surprising, as SFs represent a small wz inclusion in the zb lattice, and wz is the more stable crystal. All of the SFs have almost the same energy (certainly the difference between them is within the accuracy limit of these calculations) though the experimental observation is that the intrinsic stacking fault (SF) is the most abundant planar defect in epitaxial layers [12]. All stacking faults have a very low formation energy, implying that these defects may form very easily in both wz and zb GaN, though they are much more likely to form in the zb phase. We note that the rotation of the Ga-N bonds visible in ABF (annular bright field) STEM images of the intrinsic stacking fault[40] is reproduced in the calculated charge density of the modelling.
[…]
Wurtzite GaN
I1 I2 I3 E
[…]
N 24 33 34 25
Zincblende GaN
I E T
[…]
N 26 25 27
Accurate determination of the stacking fault thickness ($l$) is essential for calculating the band profile of stacking faults functioning as quantum wells (QWs) [6]. The calculated stacking-fault thickness $l$ for the optimized models of all the considered types of SFs are provided in Table 3. Here we use the inter-planar spacing as a measure of stacking-fault thickness, which is consistent with a recent experimental investigation based on high-resolution scanning transmission electron microscopy [12]. At 300 K, the $c$ lattice parameter of bulk wz GaN and the lattice parameter of bulk zb GaN are measured to be 5.186 Å and 4.506 Å, respectively, corresponding to inter-planar spacings of 2.593 Å and 2.602 Å in wz-GaN and zb-GaN [15, 14, 31]. Bulk zb-GaN exhibits a larger inter-planar spacing between close-packed planes compared to bulk wz-GaN, which may be attributed to stronger Coulombic interactions between third-nearest-neighbor Ga and N atoms in the wurtzite phase. Our calculated lattice parameters for the bulk wz GaN are $a$ = 3.235 Å and $c$ = 5.284 Å and for bulk zb GaN, it is $a$ = 4.578 Å. The resulting inter-planar spacing d(0001) in bulk wz-GaN is 2.642 Å and d(111) in bulk zb-GaN is 2.644 Å. These show good agreement with the experimental measurements[15, 14, 31] though perhaps the difference between the phases is a little small. SFs can be seen as small insertions into the crystal of the other crystal type, and their equilibrium inter-planar spacing may, thus, be similar to the inter-planar spacing in the non-defective materials. In Table 3, the calculated largest ($l_{max}$) and smallest ($l_{min}$) layer spacings of the SFs are presented. Our calculated stacking faults layer spacings range from 2.639 to 2.643 Å for wz GaN and 2.643 to 2.644 Å for zb GaN, showing a very small variation; plots of the evolution of the layer spacing along the $c$ axis of the simulation cell are given in the supporting information. (The alternative approach to define SF thickness using multiples of the Ga-N bond length [4, 20] gives similar insight but with less detail, so we do not present it.)
Table 2: Calculated range of layer spacings in stacking faults with respect to ideal crystals of wz and zb GaN.
Stacking faults based on Wurtzite GaN
Bulk I1 I2 I3 E
2.642 2.640–2.643 2.640–2.643 2.639–2.643 2.640–2.642
Stacking faults based on Zincblende GaN
Bulk I E T
2.644 2.643–2.644 2.643–2.644 2.643–2.644
IV Electronic structure
It is important to characterise the effect of the stacking faults on the electronic structure of the materials (both wz and zb GaN). While SFs can be seen as simply a change in the layer ordering, they can also be seen as a small inclusion of the other crystal (i.e. zb in wz and vice versa) which will involve a change from polar to non-polar (or vice versa) material, with potentially significant effects on the electronic structure.
In the first instance, we plot band-resolved charge densities to visualise the changes around the Fermi level in real-space. In the usual way, we define the density for band $n$ as $\rho_{n}(\mathbf{r})=\sum_{\mathbf{k}}w_{\mathbf{k}}|\psi_{n\mathbf{k}}(\mathbf{r})|^{2}$ where $w_{\mathbf{k}}$ is the weight for each point in the Brillouin zone; the resulting plots are a solid state equivalent of a molecular orbital plot. These band densities show similar trends for all SFs in each material, so we show here the bands for the most common SF in wz (I1) in Fig. 3 and in zb (I) in Fig. 4. Band densities for all other SFs are given in the supporting information.
In both wz and zb the bands near the top of the valence band (Fig. 3(a)-(d) in wz and Fig. 4(a)-(d) in zb) are localised on one side of the SF, and the bands near the bottom of the conduction band (Fig. 3(e)-(h) in wz and Fig. 4(e)-(h) in zb) are localised on the other side (though this is perhaps a little clearer in zb). Similar behaviour is seen in all other stacking faults in both materials.
Figure 3: Band-resolved densities of Intrinsic-1 stacking fault of wz GaN, here shown with six layers between stacking faults (/3AB/6CB/3AB/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. Note that there are two stacking faults in the simulation cell, 25% and 75% of the way along the cell.
However, the localisation occurs on different sides of the SF in wz and in zb: in wz, the valence bands are localised on the Ga-polar side of the stacking fault (i.e. on the Ga side of the Ga-N bonds perpendicular to the SF, or below the stacking fault as plotted) with the conduction bands localised on the N-polar side. By contrast, in zb the opposite trend is seen, with the valence bands on the N-polar side of the SF (above the stacking fault as plotted) and conduction bands on the Ga-polar side (below the stacking fault as plotted).
Figure 4: Band-resolved densities of Intrinsic stacking fault for zb GaN along (111), here shown with eight layers between stacking faults (/4ABC/BC/4ABC/). The bands below the Fermi energy are shown in panels (a-d), and the bands above the Fermi energy are shown in panels (e-h). The isosurfaces were plotted at a density of 0.0005 electrons per Bohr3. There is one stacking fault, halfway along the simulation cell.
We seek to identify the mechanism underlying this localisation of the bands on one side of the SF, and the origin of the difference between the different behaviour in the two materials. We characterise the problem by comparing the electrostatic potential in the faulted material with perfect material, and consider the potential along the $c$-axis of the simulation cell, taking a planar average in the aba-b plane. The resulting potential trace is shown for the most stable SFs in the top panels of Fig. 5(a) and (b) for wz and zb respectively (we again show plots for all other SFs in the supplementary information).
We see that the potential changes rapidly across the SFs in both materials (noting that there are two SFs in the wz simulation cell) with periodicity leading to a slow variation of the potential across the rest of the simulation cell. We can get insight into the extent of charge localisation by plotting the planar average of the density difference between the faulted and perfect cells (middle panel) and by integrating this along the c axis (bottom panel). We can see the expected consistency between charge difference and potential difference.
Figure 5: Potential difference between faulted and perfect material, averaged in the x-y plane (top), density difference found in the same way (middle) and density difference integrated along z (bottom) for (a) I1 SF in wz and (b) Intrinsic SF in zb.
It is clear from Fig. 5(a) that, relative to perfect material, charge accumulates on the Ga side of the SF in wz in agreement with Fig. 3. In Fig. 5(b) we see that the opposite is true in zb, with charge reducing on the Ga side of the SF and accumulating on the N side, consistent with both Fig. 4 and the change in the potential. The sign difference between the two cases occurs because the effect of a zb inclusion in wz is the exact opposite of a wz inclusion in zb. (In terms of Fig. 5(a), for the potential we plot $V_{Faulted}-V_{wz}$ which in the area around the SF will be $V_{zb}-V_{wz}$; in Fig. 5(b) in the area around the SF we will effectively be plotting $V_{wz}-V_{zb}$.) The localisation of the band densities is then easily understood in terms of the potential change across the stacking faults.
Figure 6: Band edges: (a) WZ CBM; (b) ZB CBM; (c) WZ VBM; (d) ZB VBM.
It is important to understand the nature of the band alignment around the SFs (whether a type I or type II interface) as it can affect the trapping of electrons and holes and hence the optoelectronic properties[9]. We compare the band edges in simulation cells with stacking faults to cells of perfect bulk, ensuring that the deep-lying non-dispersive $d$-bands are aligned as explained in Sec. II. The band edges are shown in Fig. 6 for both wz and zb. It should be noted that, as is standard for most DFT calculations, our calculated band gaps 1.65 eV (bulk wz GaN) and 1.52 eV (bulk zb GaN) are significantly underestimated compared to the experimental values (3.51 eV and 3.30 eV for bulk wz and zb GaN respectively[38]). This underestimate is a well-known feature of standard DFT functionals, and our calculations are in good agreement with other DFT values[4].
We see that, for all SFs in zb, these defects are type II interfaces, with both valence band maximum and conduction band minimum shifted higher than in the bulk. For wz the same is true, though for the I1 SF (the most common) the conduction band minimum is extremely close to the energy in the bulk; the behavior is likely to be very sensitive to local fields and perturbations of the electronic structure, so that hints of type I behavior might be seen. This is in broad agreement with existing simulation results[20, 32, 23, 33], though we are not aware of systematic testing of the size of simulation cell in other theory work. We also see that the gap is reduced by $\sim$0.01 eV in the wz I1 SF and by $\sim$0.04 eV in zb I SF. This result is in contradiction to simple models of SFs in zb, assumed to be wz insertions into the zb matrix, which would lead to an increase of the gap. This highlights the importance of detailed electronic structure modelling of realistic models of materials. Experimental CL and PL results show bands associated with SFs in wz samples which are at lower energies than the band edge[22, 17], with shifts of $\sim$0.06–0.18 eV.
We can further investigate the electronic properties of the stacking faults by resolving them spatially via the projected density of states (pDOS), and using this data to track the local positions of the VBM and CBM within the simulation cell. Since our focus is on the band edges, the pDOS is not shown in the main text, but is provided for all the SFs in both wz and zb in the supplementary information. In general the valence band maximum is more strongly affected than the conduction band minimum. The variation of the local band gap through the simulation cell is shown in Fig. 7 for the two most common stacking faults in wz and zb, with data for all the other SFs given in the supplementary information. We see that the change in the gap is anti-correlated with the potential differences in Fig. 5, with the gap widening where the potential is negative. This also coincides with the location of the depletion in the valence bands, consistent with the shift of the VBM to lower energies. Calculated band gap changes $\Delta E_{g}$ (eV), band offsets $\Delta E_{C}$ (eV) and $\Delta E_{V}$ (eV) for the stacking faults with respect to respective bulk crystals of wz and zb GaN are presented in supporting information.
Figure 7: Local band gap plotted across simulation cell for I1 SF in wz GaN (a) and I SF in zb GaN (b).
V Conclusions
We have performed a systematic study of stacking faults in wurtzite and zincblende phases of GaN using periodic DFT calculations. We have investigated all possible types of stacking faults structures, considering their stability and electronic properties to understand the structure properties relationship. This investigation is of central importance since stacking faults are very common defects in zb III-nitride materials, and non-polar oriented wz materials. From the structural relaxation, we can observe that the stacking faults have a rather small influence on lattice parameters and atomic positions; the inter-planar distance varies very slightly around the stacking faults.
The formation energies are quite low. In wz, the I1 intrinsic stacking fault has the lowest formation energy, while all stacking faults in zb have similar energies. The formation energy in zb is much smaller than that in wz, so stacking faults are more likely to be formed in zb phase compared to the wz phase.
The analysis of the band resolved density reveals that near the Fermi level the valence bands are concentrated on the opposite side of the stacking fault to the conduction bands. However, the concentration happens on opposite sides in wz and zb, which is understood from the fact that SFs can be seen as an inclusion of the other phase.
Finally, a detailed analysis of the electronic structure shows that, in contrast to standard models, the band gap of stacking faults is smaller than that of the ideal wurtzite and zincblende phases, with a positive offset. This change in the band edges is equivalent to a type II interface in all cases, though the most common wz stacking fault is highly sensitive to local conditions.
The data that support the findings of this article are openly available[39].
Acknowledgements.
This research was funded through UKRI/EPSRC grant number EP/Y00423X/1. We gratefully acknowledge many useful discussions and insights from Professor Rachel Oliver, Dr Martin Frentrup, Dr Petr Vacek and Professor David Wallis of Cambridge University. This research was partly carried out using the high performance computing cluster in the London Centre for Nanotechnology. The authors acknowledge the use of the UCL Kathleen High Performance Computing Facility (KathleenUCL), and associated support services, in the completion of this work. We are grateful for computational support from the UK Materials and Molecular Modelling Hub, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1), for which access was obtained via the UKCP consortium and funded by EPSRC grant ref EP/P022561/1. Supporting Information
Appendix A Basis sets
The pseudo atomic orbital (PAO) basis sets[8] used for Ga and N were double-zeta plus polarization (DZP) basis sets, generated with the equal radius setting in Conquest. The gallium pseudopotential includes the $3d$ electrons as semi-core states, with the radii of the PAOs in Bohr as follows: $3d$: 3.58 $a_{0}$; $4s$ and $4p$: 8.06 and 4.33 $a_{0}$; 4d (polarization): 8.06 $a_{0}$. For N the radii are: $2s$ and $2p$: 5.48 and 2.81 $a_{0}$; 3d (polarization): 5.48 $a_{0}$. For indium, the pseudopotential includes the $4d$ electrons as semi-core states, with the radii of the PAOs in Bohr as follows: $4d$: 4.11 $a_{0}$; $5s$ and $5p$: 8.41 and 4.63 $a_{0}$; 5d (polarization): 8.41 $a_{0}$. For aluminium, the radii of the PAOs in Bohr are as follows: $3s$ and $3p$: 8.24 and 4.48 $a_{0}$; 3d (polarization): 8.24 $a_{0}$.
Appendix B Physical parameters
Here we give lattice constants of the bulk, along with convergence tests for simulation cell size.
Table 3: Comparison of calculated hexagonal lattice parameters a, c (Å) of wz and cubic lattice parameter a (Å) in zb phases of GaN, InN and AlN with available literature data.
WZ
$a$ $c$
This Study Previous Study [16] Exp. [14, 31] This Study Previous Study [16] Exp. [15, 14, 31]
GaN 3.24 3.24 3.19 5.29 5.23 5.19
InN 3.60 3.61 3.56 5.85 5.83 5.70
AlN 3.13 3.13 3.11 5.08 5.02 4.98
Zb
$a$
This Study Previous Study [11] Exp. [15, 14]
GaN 4.58 4.55 4.49
InN 5.08 5.06 4.98
AlN 4.42 4.40 4.37
Table 4: Considered models for stacking faults, number of atoms and c parameter (Å)
Stacking faults based on Wurtzite GaN
I1 Atoms $c$ I2 Atoms $c$ I3 Atoms $c$ Extrinsic Atoms $c$
AB-2CB-AB 32 21.154 AB-2CA-C-AB 36 23.798 AB-C-2BA-C-AB 40 26.441 2AB-C-2AB 36 23.798
2AB-4CB-2AB 64 42.311 2AB-4CA-C-2AB 68 44.945 2AB-C-4BA-C-2AB 72 47.598 3AB-C-3AB 52 34.370
3AB-6CB-3AB 96 63.469 3AB-6CA-C-3AB 100 66.096 3AB-C-6BA-C-3AB 104 68.741 4AB-C-4AB 68 84.936
4AB-8CB-4AB 128 84.612 4AB-8CA-C-4AB 132 87.269 4AB-C-8BA-C-4AB 136 89.891 5AB-C-5AB 84 55.521
5AB-10CB-5AB 160 105.767 5AB-10CA-C-5AB 164 108.399 5AB-C-10BA-C-5AB 168 111.039 6AB-C-6AB 100 66.102
6AB-12CB-6AB 192 126.946 6AB-12CA-C-6AB 196 129.551 6AB-C-12BA-C-6AB 200 132.198 7AB-C-7AB 116 76.677
7AB-14CB-7AB 224 279.823 7AB-14CA-C-7AB 228 284.805 7AB-C-14BA-C-7AB 232 289.773 8AB-C-8AB 132 164.729
Stacking faults based on Zincblende GaN
Intrinsic Atoms $c$ Extrinsic Atoms $c$ Twin Atoms $c$
ABC-BC-ABC 32 39.971 ABC-B-ABC 28 34.975 ABC-BA-CABC 36 44.968
2ABC-BC-ABC 44 54.961 2ABC-B-ABC 40 49.964 2ABC-BA-C2ABC 60 74.946
2ABC-BC-2ABC 56 69.950 2ABC-B-2ABC 52 64.953 3ABC-BA-C3ABC 84 104.925
3ABC-BC-2ABC 68 84.939 3ABC-B-2ABC 64 79.943 4ABC-BA-C4ABC 108 134.903
3ABC-BC-3ABC 80 99.928 3ABC-B-3ABC 76 94.932 5ABC-BA-C5ABC 132 164.882
4ABC-BC-3ABC 92 114.917 4ABC-B-3ABC 88 109.921 6ABC-BA-C6ABC 156 194.859
4ABC-BC-4ABC 104 129.906 4ABC-B-4ABC 100 124.910 7ABC-BA-C7ABC 180 224.838
5ABC-BC-4ABC 116 144.896 5ABC-B-4ABC 112 139.899
5ABC-BC-5ABC 128 159.885 5ABC-B-5ABC 124 154.889
Table 5: Calculated values of formation energy (mJ/m2) of stacking faults with respect to increasing number of layers
Stacking faults based on Wurtzite GaN
[…]
AB-2CB-AB 18.627 AB-2CA-C-AB 26.806 AB-C-2BA-C-AB 34.301 2AB-C-2AB 54.989
2AB-4CB-2AB 17.058 2AB-4CA-C-2AB 26.908 2AB-C-4BA-C-2AB 34.463 3AB-C-3AB 54.711
3AB-6CB-3AB 17.051 3AB-6CA-C-3AB 27.275 3AB-C-6BA-C-3AB 35.346 4AB-C-4AB 54.643
4AB-8CB-4AB 16.709 4AB-8CA-C-4AB 27.990 4AB-C-8BA-C-4AB 35.592 5AB-C-5AB 54.918
5AB-10CB-5AB 16.546 5AB-10CA-C-5AB 27.935 5AB-C-10BA-C-5AB 35.898 6AB-C-6AB 55.270
6AB-12CB-6AB 16.936 6AB-12CA-C-6AB 28.162 6AB-C-12BA-C-6AB 36.349 7AB-C-7AB 55.357
7AB-14CB-7AB 17.123 7AB-14CA-C-7AB 28.199 7AB-C-14BA-C-7AB 36.331 8AB-C-8AB 55.426
Stacking faults based on Zincblende GaN
[…]
ABC-BC-ABC -32.267 ABC-B-ABC -31.159 ABC-BA-CABC -34.786
2ABC-BC-ABC -33.466 2ABC-B-ABC -35.127 2ABC-BA-C2ABC -36.002
2ABC-BC-2ABC -33.567 2ABC-B-2ABC -35.010 3ABC-BA-C3ABC -34.804
3ABC-BC-2ABC -33.487 3ABC-B-2ABC -36.281 4ABC-BA-C4ABC -34.467
3ABC-BC-3ABC -33.470 3ABC-B-3ABC -34.744 5ABC-BA-C5ABC -34.190
4ABC-BC-3ABC -33.150 4ABC-B-3ABC -34.604 6ABC-BA-C6ABC -33.826
4ABC-BC-4ABC -32.934 4ABC-B-4ABC -34.413 7ABC-BA-C7ABC -33.494
5ABC-BC-4ABC -32.783 5ABC-B-4ABC -34.217
5ABC-BC-5ABC -34.413 5ABC-B-5ABC -33.901
Figure 8: Evolution of layer spacing along the c-axis of the simulation cell of the wz GaN SFs with respect to calculated bulk wz GaN. I1 (a), I2 (b), I3 (c) and Extrinsic (d) SFs
Figure 9: Evolution of layer spacing along the c-axis of the simulation cell of the ZB GaN SFs with respect to calculated bulk zb GaN. Intrinsic (a), Extrinsic (b) and Twin (c)
Appendix C Band Densities
The band densities for the I1 stacking fault in wz GaN and the I stacking fault in zb GaN are shown in the main text. Here we show the other band densities in Figs. 10 to 14.
Figure 10: Band Density of Intrinsic-2 (4AB-8CA-C-4AB) stacking fault of wz GaN. (a) Band 591, (b) Band 592, (c) Band 593, (d) Band 594, (e) Band 595, (f) Band 596, (g) Band 597, (h) Band 598. Fermi Energy (Ef: -5.86 eV) lies in between band 594 and band 595.
Figure 11: Band Density of Intrinsic-3 (4AB-C-8BA-C-4AB) stacking fault of wz GaN. (a) Band 609, (b) Band 610, (c) Band 611, (d) Band 612, (e) Band 613, (f) Band 614, (g) Band 615, (h) Band 616. Fermi Energy (Ef: -5.86 eV) lies in between band 612 and band 613.
Figure 12: Band Density of Extrinsic (6AB-C-6AB) stacking fault of wz GaN. (a) Band 447 , (b) Band 448 , (c) Band 449, (d) Band 450, (e) Band 451 , (f) Band 452 , (g) Band 453 , (h) Band 454. Fermi Energy (Ef: -5.80 eV) lies in between band 450 and band 451.
Figure 13: Band Density of Extrinsic (4ABC-B-4ABC) stacking fault for zb GaN. (a) Band 447 , (b) Band 448 , (c) Band 449, (d) Band 450, (e) Band 451 , (f) Band 452 , (g) Band 453 , (h) Band 454. Fermi Energy (Ef = -0.2146 Ha) lies in between band 450 and band 451.
Figure 14: Band density of Twin (4ABC-BAC-4ABC) stacking fault for zb GaN. (a) Band 483 , (b) Band 484 , (c) Band 485, (d) Band 486, (e) Band 487 , (f) Band 488 , (g) Band 489 , (h) Band 490. Fermi Energy (Ef = –0.2145805742 Ha ) lies in between band 486 and band 487.
Appendix D Potentials
Figure 15: Average potential of stacking faults based on wz GaN (a) I1 (3AB-6CB-3AB) (b) I2 (4AB-8CA-C-4AB) (c) I3 (4AB-C-8BA-C-4AB) (d) Extrinsic (6AB-C-6AB).
Figure 16: Average potential of stacking faults based on zb GaN (a) intrinsic (b) extrinsic and (c) twin.
Appendix E Projected DOS
Figure 17: Projected density of states (pDOS) for wurtzite bulk and atoms located at the SFs, showing total (blue) and decomposed by angular momentum: (a) and (b) bulk (Ga and N respectively); (c) and (d) I1 SF (Ga and N respectively); (e) and (f) I2 SF (Ga and N respectively); (g) and (h) I3 SF (Ga and N respectively); (i) and (j) Extrinsic SF (Ga and N respectively).
Figure 18: Projected density of states (pDOS) for zincblende bulk and atoms located at the SFs, showing total (blue) and decomposed by angular momentum: (a) and (b) bulk (Ga and N respectively); (c) and (d) intrinsic SF (Ga and N respectively); (e) and (f) extrinsic SF (Ga and N respectively); (g) and (h) twin SF (Ga and N respectively).
Appendix F Band Gap Changes
Table 6: Calculated values of band gap difference $\Delta E_{g}$ (eV), band offsets $\Delta E_{C}$ (eV) and $\Delta E_{V}$ (eV) for the stacking faults with respect to ideal crystals of wz and zb GaN
Stacking faults based on Wurtzite GaN
Intrinsic 1 (I1) Intrinsic 2 (I2) Intrinsic 3 (I3) Extrinsic
$\Delta E_{g}$ -0.015 (-0.050[4]) -0.046 (-0.120[4]) -0.043 -0.095 (-0.140[4])
$\Delta E_{C}$ 0.005 (-0.045[4]) 0.039 (-0.096[4]) 0.028 0.054(-0.160[4])
$\Delta E_{V}$ 0.021(0.011[4]) 0.085 (0.019[4]) 0.071 0.149(0.028[4])
Stacking faults based on Zincblende GaN
Intrinsic Extrinsic Twin
$\Delta E_{g}$ -0.037 -0.032 -0.033
$\Delta E_{C}$ 0.064 0.060 0.059
$\Delta E_{V}$ 0.101 0.092 0.092
Figure 19: Band gap evolution along c-axis of simulation cell for stacking faults in wz GaN. I1 (a), I2 (b), I3 (c) and Extrinsic (d)
Figure 20: Band gap evolution along c-axis of simulation cell for stacking faults of zb GaN. Intrinsic (a), Extrinsic (b) and Twin (c).